Tension is maintained in a string of 29.0 N and transverse wave is observed to pass with a speed of 27.0 m/s. (a) What linear mass density this string has? (b) If the tension is reduced half of its initial value, what speed will a transverse wave can be propagated through the string?


1
Expert's answer
2022-05-02T13:41:40-0400

(a)

"v=\sqrt{T/\mu}"

"\mu=T/v^2=\frac{29.0\:\rm N}{(27.0\:\rm m/s)^2}=0.040\rm\:kg/m"

(b)

"v'=v/\sqrt{2}=27.0/\sqrt{2}=19.1\:\rm m/s"